# Probing Planck scale physics with Belle II and LHCb

### v1

@@ 2.1 Time evolution of single $B$ mesons
$E_{0}$ $=$ $\left|0\right\rangle\left\langle 0\right|\,,$ (6)
$E_{1}$ $=$ $\frac{1}{2}\left[e^{-\left(2im_{L}+\Gamma_{L}+\lambda\right)t/2}+e^{-\left(2im_{H}+\Gamma_{H}+\lambda\right)t/2}\right]\left(\left|B^{0}\right\rangle\left\langle B^{0}\right|+\left|\bar{B^{0}}\right\rangle\left\langle\bar{B^{0}}\right|\right)$ (7)
@@
$E_{2}$ $=$ $\sqrt{\frac{{\rm Re}[(p-q)/(p+q)]}{|p|^{2}-|q|^{2}}\left(1-e^{-\Gamma_{L}\,t}-\left(|p|^{2}-|q|^{2}\right)^{2}\frac{\left|1-e^{-(\Gamma+\lambda-i\Delta m)t}\right|^{2}}{(1-e^{-\Gamma_{H}\,t})}\right)}$ (8)
@@
$E_{3}$ $=$ $\sqrt{\frac{{\rm Re}[(p-q)/(p+q)]}{(|p|^{2}-|q|^{2})(1-e^{-\Gamma_{H}\,t})}}\Big{(}\frac{1-e^{-\Gamma_{H}\,t}+\left(1-e^{-(\Gamma+\lambda-i\Delta m)t}\right)\left(|p|^{2}-|q|^{2}\right)}{2p}(p+q)\left|0\right\rangle\left\langle B^{0}\right|$ (9)
@@
$E_{4}$ $=$ $\frac{1}{2}e^{-\Gamma_{L}t/2}\sqrt{1-e^{-\lambda t}}\left(\left|B^{0}\right\rangle\left\langle B^{0}\right|+\left|\bar{B^{0}}\right\rangle\left\langle\bar{B^{0}}\right|+\frac{p}{q}\left|B^{0}\right\rangle\left\langle\bar{B^{0}}\right|+\frac{q}{p}\left|\bar{B^{0}}\right\rangle\left\langle B^{0}\right|\right)\,,$ (10)
$E_{5}$ $=$ $\frac{1}{2}e^{-\Gamma_{H}t/2}\sqrt{1-e^{-\lambda t}}\left(\left|B^{0}\right\rangle\left\langle B^{0}\right|+\left|\bar{B^{0}}\right\rangle\left\langle\bar{B^{0}}\right|-\frac{p}{q}\left|B^{0}\right\rangle\left\langle\bar{B^{0}}\right|-\frac{q}{p}\left|\bar{B^{0}}\right\rangle\left\langle B^{0}\right|\right)\,.$ (11)
@@ 2.2 Time evolution of correlated $B$ mesons
where $\rho_{1}(t)=\rho_{B^{0}}(t)$, $\rho_{2}(t)=\rho_{\bar{B^{0}}}(t)$ which are given in eq. (20), while $\rho_{3/4}(t)=\sum_{i}K_{i}\rho_{3/4}(0)K_{i}^{\dagger}$, where
@@ 3 Meson anti-meson mixing
The current world average for $\Delta m_{d}$, as cited in the Particle Data Group (PDG), stands at $(0.5065\pm 0.0019),\rm ps^{-1}$ pdg . This value represents the amalgamation of measurements of $\Delta m_{d}$ derived from a spectrum of experiments, including OPAL opal , ALEPH aleph , DELPHI delphi , L3 l3 , CDF cdf , BaBar babar , Belle belle , D0 do , and LHCb LHCb:2011vae ; LHCb:2012mhu ; LHCb:2013fep . Consequently, it becomes apparent that the process of determining $\Delta m_{d}$ in experiments like LHCb, CDF, D0, Belle, and BaBar can be affected by the presence of the decoherence parameter $\lambda$. Therefore, reevaluating the determination of $\Delta m_{d}$ in these experiments while considering decoherence can provide constraints on the $\lambda$ parameter. On the other hand, the LEP experiments gauge $\Delta m_{d}$ through time-independent measurements. In the following section, we will demonstrate that these measurements are likewise affected by the presence of decoherence.
@@ 5 Purely leptonic decays of $B$ mesons
$N_{CP}^{LR}(t)$ $=$ $\frac{1}{2}e^{-\Gamma t}\Bigg{[}\left(|\lambda_{L}|^{2}-|\lambda_{R}|^{-2}\right)\cosh\left(\frac{\Delta\Gamma t}{2}\right)-\left(|\lambda_{L}|^{2}-|\lambda_{R}|^{-2}\right)e^{-\lambda t}\cos\left(\Delta mt\right)$
@@
$D_{CP}^{LR}$ $=$ $\frac{\left\{2+\left(|\lambda_{L}|^{2}+|\lambda_{R}|^{-2}\right)\right\}}{1-y^{2}}+\frac{(1+\lambda^{\prime})\left\{2-\left(|\lambda_{L}|^{2}+|\lambda_{R}|^{-2}\right)\right\}}{(1+\lambda^{\prime})^{2}+x^{2}}+\frac{4\,y\,{\rm Re}\left(\lambda_{L}+\lambda_{R}^{-1}\right)}{1-y^{2}}$ (77)
$-\frac{4\,x\,{\rm Im}\left(\lambda_{L}+\lambda_{R}^{-1}\right)}{(1+\lambda^{\prime})^{2}+x^{2}}\,.$
@@ 6 Nonleptonic decays of $B$ mesons
The time-dependent $CP$ asymmetry as given in eq. (113) with $\lambda=0$ is used to determine quantities like $\sin 2\beta$ and $\sin 2\beta_{s}$ where $\beta\equiv\arg\left[-(V_{cd}V_{cb}^{*})/(V_{td}V_{tb}^{*})\right]$ and $\beta_{s}\equiv\arg\left[-(V_{ts}V_{tb}^{*})/(V_{cs}V_{cb}^{*})\right]$ are the angles of the unitarity triangle. For e.g., the most statistically precise measurement of $\sin 2\beta$ is obtained through a time-dependent angular analysis of $B^{0}_{d}\to J/\psi\,K_{S}$ which is also regarded as the golden channel to determine $\sin 2\beta$. For this decay channel, $\Delta\Gamma=0$, $\lambda_{f}=1$ and in the limit of neglecting the penguin loop contributions to $b\to c\bar{c}s$ decay, we get

### v2

# Probing quantum decoherence at Belle II and LHCb

@@ 2.1 Time evolution of single $B$ mesons
$E_{0}(t)$ $=$ $\left|0\right\rangle\left\langle 0\right|\,,$ (6)
$E_{1}(t)$ $=$ $\frac{1}{2}\left[e^{-\left(2im_{L}+\Gamma_{L}+\lambda\right)t/2}+e^{-\left(2im_{H}+\Gamma_{H}+\lambda\right)t/2}\right]\left(\left|B^{0}\right\rangle\left\langle B^{0}\right|+\left|\bar{B^{0}}\right\rangle\left\langle\bar{B^{0}}\right|\right)$ (7)
@@
$E_{2}(t)$ $=$ $\sqrt{\frac{{\rm Re}[(p-q)/(p+q)]}{|p|^{2}-|q|^{2}}\left(1-e^{-\Gamma_{L}\,t}-\left(|p|^{2}-|q|^{2}\right)^{2}\frac{\left|1-e^{-(\Gamma+\lambda-i\Delta m)t}\right|^{2}}{(1-e^{-\Gamma_{H}\,t})}\right)}$ (8)
@@
$E_{3}(t)$ $=$ $\sqrt{\frac{{\rm Re}[(p-q)/(p+q)]}{(|p|^{2}-|q|^{2})(1-e^{-\Gamma_{H}\,t})}}\Big{(}\frac{1-e^{-\Gamma_{H}\,t}+\left(1-e^{-(\Gamma+\lambda-i\Delta m)t}\right)\left(|p|^{2}-|q|^{2}\right)}{2p}(p+q)\left|0\right\rangle\left\langle B^{0}\right|$ (9)
@@
$E_{4}(t)$ $=$ $\frac{1}{2}e^{-\Gamma_{L}t/2}\sqrt{1-e^{-\lambda t}}\left(\left|B^{0}\right\rangle\left\langle B^{0}\right|+\left|\bar{B^{0}}\right\rangle\left\langle\bar{B^{0}}\right|+\frac{p}{q}\left|B^{0}\right\rangle\left\langle\bar{B^{0}}\right|+\frac{q}{p}\left|\bar{B^{0}}\right\rangle\left\langle B^{0}\right|\right)\,,$ (10)
$E_{5}(t)$ $=$ $\frac{1}{2}e^{-\Gamma_{H}t/2}\sqrt{1-e^{-\lambda t}}\left(\left|B^{0}\right\rangle\left\langle B^{0}\right|+\left|\bar{B^{0}}\right\rangle\left\langle\bar{B^{0}}\right|-\frac{p}{q}\left|B^{0}\right\rangle\left\langle\bar{B^{0}}\right|-\frac{q}{p}\left|\bar{B^{0}}\right\rangle\left\langle B^{0}\right|\right)\,.$ (11)
@@
2.1.1 Theoretical limit on the decoherence parameter $\lambda$
For the time evolution of the density operator to be completely positive, the following inequality must hold for any $t\geq 0$ Caban:2005ue :
$\delta^{2}_{L}\left(1-2e^{-(\Gamma+\lambda)t}\cos(\Delta mt)+e^{-2(\Gamma+\lambda)t}\right)\leq\left(1-e^{-\Gamma_{H}t}\right)\left(1-e^{-\Gamma_{L}t}\right)\,.$ (21)
Here $\delta_{L}=|p|^{2}-|q|^{2}$. Given that the parameters $\Gamma_{L}$, $\Gamma_{H}$, $\Delta m$, and $\delta_{L}$ are known, the above inequality can be utilized to obtain an upper bound on the decoherence parameter $\lambda$. This bound serves as a theoretical upper limit, and any experimentally determined limit on $\lambda$ should not exceed this bound. Hence, this inequality also imposes constraints on the validity of the model under consideration.
For $t\geq 0$ and $\lambda\geq 0$, the above inequality can be written in the following form Caban:2005ue
$0\leq\lambda\leq\lambda_{\rm max}\simeq\frac{1}{\delta_{L}}\sqrt{\Gamma_{L}\Gamma_{H}-\delta^{2}_{L}(\Delta m)^{2}}-\Gamma\,.$ (22)
For the neutral $B_{d}$ meson system, $\Delta m_{B_{d}}=(0.5065\pm 0.0019)\,{\rm ps^{-1}}$, $\tau_{B_{d}}=1/\Gamma_{B_{d}}=(1.519\pm 0.004)\,{\rm ps}$ and $\delta_{L}=(-0.5\pm 0.4)\times 10^{-3}$. These values are taken from Particle Data Group (PDG) pdg . As for this system $\Delta\Gamma=\Gamma_{L}-\Gamma_{H}\approx 0$, we assume $\Gamma_{L}\approx\Gamma_{H}\approx\Gamma$. Substituting these values in eq. (22), we get $\lambda_{\rm max}=7.9\times 10^{14}\,{\rm s^{-1}}$ at 95% C.L. The current experimental upper limit on $\lambda$ is $3\times 10^{10}$ $\rm s^{-1}$ at 95% C.L. Alok:2015iua which is well within the allowed limit.
For the neutral $B_{s}$ meson system, $\Delta m_{B_{s}}=(17.765\pm 0.006)\,{\rm ps^{-1}}$, $\Gamma_{B_{s}}=(65.73\pm 0.23)\times 10^{10}\,{\rm s^{-1}}$, $\tau_{B_{sL}}=1/\Gamma_{B_{sL}}=(1.431\pm 0.007)\,{\rm ps}$, $\tau_{B_{sH}}=1/\Gamma_{B_{sH}}=(1.624\pm 0.009)\,{\rm ps}$ and $\delta_{L}=(-0.15\pm 0.70)\times 10^{-3}$. By substituting these values, taken from PDG pdg , into eq. (22), we derive $\lambda_{\rm max}=3.64\times 10^{16}\,{\rm s^{-1}}$ at 95% C.L. Currently, there are no experimental limits available for the decoherence parameter in the $B_{s}$ meson system. However, like the $B_{d}$ meson system it is expected that the experimental limit on $\lambda$ in $B_{s}$ system should also fall within the acceptable range.
@@ 2.2 Time evolution of correlated $B$ mesons
where $\rho_{1}(t)=\rho_{B^{0}}(t)$, $\rho_{2}(t)=\rho_{\bar{B^{0}}}(t)$ which are given in eq. (20), while $\rho_{3/4}(t)=\sum_{i}E_{i}(t)\rho_{3/4}(0)E_{i}^{\dagger}(t)$, where
@@ 3 Meson anti-meson mixing
The current world average for $\Delta m_{d}$, as cited in the PDG, stands at $(0.5065\pm 0.0019),\rm ps^{-1}$ pdg . This value represents the amalgamation of measurements of $\Delta m_{d}$ derived from a spectrum of experiments, including OPAL opal , ALEPH aleph , DELPHI delphi , L3 l3 , CDF cdf , BaBar babar , Belle belle , D0 do , and LHCb LHCb:2011vae ; LHCb:2012mhu ; LHCb:2013fep . Consequently, it becomes apparent that the process of determining $\Delta m_{d}$ in experiments like LHCb, CDF, D0, Belle, and BaBar can be affected by the presence of the decoherence parameter $\lambda$. Therefore, reevaluating the determination of $\Delta m_{d}$ in these experiments while considering decoherence can provide constraints on the $\lambda$ parameter. On the other hand, the LEP experiments gauge $\Delta m_{d}$ through time-independent measurements. In the following section, we will demonstrate that these measurements are likewise affected by the presence of decoherence.
@@ 5 Purely leptonic decays of $B$ mesons
$N_{CP}^{LR}(t)$ $=$ $\frac{1}{2}e^{-\Gamma t}|A_{L}|^{2}\Bigg{[}\left(|\lambda_{L}|^{2}-|\lambda_{R}|^{-2}\right)\cosh\left(\frac{\Delta\Gamma t}{2}\right)-\left(|\lambda_{L}|^{2}-|\lambda_{R}|^{-2}\right)e^{-\lambda t}\cos\left(\Delta mt\right)$
@@
$D_{CP}^{LR}$ $=$ $\frac{\left\{2+\left(|\lambda_{L}|^{2}+|\lambda_{R}|^{-2}\right)\right\}}{1-y^{2}}+\frac{(1+\lambda^{\prime})\left\{2-\left(|\lambda_{L}|^{2}+|\lambda_{R}|^{-2}\right)\right\}}{(1+\lambda^{\prime})^{2}+x^{2}}+\frac{2\,y\,{\rm Re}\left(\lambda_{L}+\lambda_{R}^{-1}\right)}{1-y^{2}}$ (79)
$-\frac{2\,x\,{\rm Im}\left(\lambda_{L}+\lambda_{R}^{-1}\right)}{(1+\lambda^{\prime})^{2}+x^{2}}\,.$
@@ 6 Nonleptonic decays of $B$ mesons
The time-dependent $CP$ asymmetry as given in eq. (115) with $\lambda=0$ is used to determine quantities like $\sin 2\beta$ and $\sin 2\beta_{s}$ where $\beta\equiv\arg\left[-(V^{\phantom{*}}_{cd}V_{cb}^{*})/(V^{\phantom{*}}_{td}V_{tb}^{*})\right]$ and $\beta_{s}\equiv\arg\left[-(V^{\phantom{*}}_{ts}V_{tb}^{*})/(V^{\phantom{*}}_{cs}V_{cb}^{*})\right]$ are the angles of the unitarity triangle. For e.g., the most statistically precise measurement of $\sin 2\beta$ is obtained through a time-dependent angular analysis of $B^{0}_{d}\to J/\psi\,K_{S}$ which is also regarded as the golden channel to determine $\sin 2\beta$. For this decay channel, $\Delta\Gamma=0$, $\lambda_{f}=1$ and in the limit of neglecting the penguin loop contributions to $b\to c\bar{c}s$ decay, we get